OpenStudy (anonymous):

.A study of two hundred teens found that the number of hours they spend on social networking sites each week is normally distributed with a mean of 12 hours. The population standard deviation is 4 hours. What is the margin of error for a 98% confidence interval? 0.122 0.659 0.295 0.313

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

what is your z score for 98% CI, then adjust it by the standard error:\[Z*SE=z(sd/\sqrt n)\]

OpenStudy (amistre64):

in other words:\[CI=mean\pm z(sd/\sqrt n)\] so your margin of error is the second term

OpenStudy (anonymous):

Okay, i cant remember the formula for z score.

OpenStudy (amistre64):

there is no formula, there is a messy integration, which people did long time ago and made tables from

OpenStudy (amistre64):

since we want a 98% CI about the mean, then a left tail area will give us 98 + 1, or 99% to look up in a left tail table

OpenStudy (amistre64):

or if you have a ti83, invnorm(.9900)

OpenStudy (amistre64):